# When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails

### v2

@@ A.2.2 COG Generation Parameters
Table 10: Generation Parameter Settings
Stage Parameters Values
@@
Table 11: Training Hyperparameters
@@
Finetuning Type full Learning Rate 2e-6
Cutoff Length 8192 Epochs 3.0
Batch Size 2 Warmup Ratio 0.1
Gradient Accumulation Steps 4
Table 12: Benchmark Implementation Details
Safety Benchmarks Reasoning Benchmarks
Parameters Values Parameters Values
@@ A.3.3 Benchmark Hyperparameters Details
Appendix B Additional Experimental Results
B.1 Initial Investigation by Switching Thinking Mode
Recent works  (Zhou et al., 2025a; Zhang et al., 2025) have discovered that LRMs tend to answer harmful questions. To further validate whether the thinking trajectory influences the safety of LLMs, we compare multiple LRMs with the thinking mode switch on & off. The evaluation is performed on 2k harmful queries from Wildjailbreak  (Jiang et al., 2024a), with safety assessed using Llama-Guard-3-8B (Llama Team, 2024) as the automatic judge.
The thinking trajectory substantially undermines the safety of LRMs. As shown in Table 13, LRMs become consistently safer when the thinking mode is disabled, but their safety degrades notably once it is enabled. This trend holds across different model scales: enabling the thinking trajectory systematically increases the likelihood of harmful responses. For example, the harmful answer rate of Qwen3-32B rises by 10.55% when the mode is enabled, while the 8B and 14B variants also exhibit noticeable increases (6.04% and 3.35%, respectively). These results indicate that content generated within the thinking trajectory can directly drive LRMs to produce harmful outputs.
Table 13: Percentage of harmful answers (”Answer”) and refusals (”Refuse”) on Wildjailbreak. Higher ’Answer’ indicates lower safety. $\Delta=$ Disabled - Enabled.
Model Enabled Disabled $\Delta$ (Answer)
Answer Refuse Answer Refuse
Qwen3-8B 41.40 58.60 35.36 64.64 6.04-6.04
Qwen3-14B 35.65 64.35 32.30 67.70 3.35-3.35
Qwen3-32B 41.45 58.55 30.90 79.10 10.55-10.55
B.2 Analysis: PCA Analysis of 8B and 14B Models
We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering. The observed differences are visualized in Figure 6 and 7, while quantitative results across all configurations are reported in Table 14.
Table 14: Safety distance analysis of Qwen3 models by model size. Values represent perpendicular distances from cluster centroids to the linear decision boundary. Positive $\Delta$ values indicate movement away from the boundary (greater margin, safer region), while negative $\Delta$ values indicate movement toward the boundary (reduced margin, less safe region).
Method Qwen3-8B Qwen3-14B Qwen3-32B
Harmful Reasoning Harmful Reasoning Harmful Reasoning
$\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$
Base 13.24 14.48 15.54 9.05 11.20 12.18
SafR 18.55 15.02 18.95 8.33 16.18 11.50
+5.30 +0.54 +3.41 -0.73 +4.98 -0.68
SafB 19.51 12.95 19.14 9.63 15.16 10.41
+6.27 -1.54 +3.60 +0.58 +3.96 -1.77
SafeChain 13.45 5.87 12.59 4.26 11.58 6.92
+0.21 -8.61 -2.95 -4.79 +0.38 -5.26
@@ B.2 Analysis: PCA Analysis of 8B and 14B Models
Both Safety Recomposition (SafR) and Safety Backtrack (SafB) fine-tuning strategies contribute to enhanced model safety, as evidenced by the centroid shifts observed in Figure 6 and Figure 7, and further quantified by the increase in Safety Distance (e.g., for the 8B model: from 20.05 to 25.68 after SafR). Notably, SafR yields consistently superior clustering performance, with higher Silhouette Scores than SafB across both model sizes. This suggests that SafR not only improves safety alignment but also leads to more coherent and well-separated internal representations.
Figure 6: PCA of the Qwen3-8B representation space.
Figure 7: PCA of the Qwen3-14B representation space.
B.3 Preservation of Reasoning Token Usage
Table 15 reports the average token counts for GPQA and AIME benchmarks across different training methods. SafR and SafB maintain token usage close to the Base model, while improving safety. In contrast, other baselines exhibit larger deviations, suggesting potential inefficiencies in reasoning. These results demonstrate that SafR and SafB enhance model safety without significantly affecting intrinsic reasoning behavior, selectively reinforcing safety-related features while preserving core reasoning capabilities.
Table 15: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3-8B, Qwen3-14B, and Qwen3-32B across different training methods.
Model Method GPQA-Diamond AIME
Qwen3-8B Base 7553.84 14895.49
@@ B.3 Preservation of Reasoning Token Usage
B.4 Analysis: Reasoning Safety Analysis
Prior work has primarily focused on improving the safety of reasoning through external interventions (safety for reasoning). In this section, we take the opposite view and explore whether a model’s internal reasoning ability can be leveraged to enhance overall safety (reasoning for safety). To evaluate our approach, we compare it against the recently proposed R2D method  (Zhu et al., 2025) as a representative baseline. The experimental results are shown in Table 16.
Table 16: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. Lower is better for SafB, SafR, W-JB, and JB-B; higher is better for GPQA and AIME.
Model SafB SafR W-JB JB-B GPQA AIME
@@ B.4 Analysis: Reasoning Safety Analysis
Based on the experimental results shown in the Table 16, we observe that applying reasoning-for-safety methods to LRMs indeed improves safety performance. However, similar to SafeKey and Star-1, these methods tend to sacrifice reasoning ability, failing to strike a proper balance between safety and reasoning capabilities. Although our methods partially address this issue, maintaining the integrity of reasoning ability remains a significant challenge.
Appendix C Self-Jailbreak Category Example
C.1 Benign Reframing
@@
C.2 Logical Fallacies
@@ C.3 Warning
C.4 Harm Identification
The example of Warning in Harm Identification is presented in Figure 11, where LRM answers the harmful quesiton as the regular one.
Appendix D Prompt Design
@@ D.2 Classification Prompt
D.3 Safety Recomposition Prompt
For the Safety Recomposition stage, our prompt is constructed by concatenating several components. Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts” field in a JSON file. Then, we combine the ”main_prompt”, the selected ”sub_prompt”, and a ”format_prompt” to form the final prompt. The detailed structure is as shown in Figure 15
D.4 Safety Backtrack Prompt
In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process. In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases”, selected based on the classification result. This helps guide the model more smoothly and maintain coherence in the final output. The detailed structure is as shown in Figure 16
D.5 Integration Prompt
@@
D.6 BoN Ranking Prompt
In this step, we employ a Best-of-N (BoN) ranking mechanism to select the most appropriate final output from multiple candidates generated by the model. The ranking prompt used to guide this selection process is shown in Figure 18.
@@ D.6 BoN Ranking Prompt
Figure 18: Prompt structure used in the Best-of-N (BoN) ranking stage to select the best output from multiple candidates.

### v3

@@ A.2.1 COG Generation Parameters
Stage Parameter Value
@@
Table 10: Generation parameter settings.
Parameter Value
Finetuning Type Full
Learning Rate 2e-6
Cutoff Length 8192
Epochs 3.0
Batch Size 2
Warmup Ratio 0.1
Gradient Accumulation Steps 4
Table 11: Training hyperparameters.
Safety Benchmarks Reasoning Benchmarks
@@
MATH500 rollout 1
HumanEval rollout 1
Table 12: Benchmark implementation details.
@@ A.3.3 Benchmark Hyperparameters Details
Appendix B Initial Investigation by Switching Thinking Mode
Recent works (Zhou et al., 2025a; Zhang et al., 2025b) have observed that LRMs tend to answer harmful questions. To further examine whether the thinking trajectory causally influences model safety, we evaluate multiple LRMs with the thinking mode switched on and off. Specifically, we test 2k harmful queries from WildJailbreak (Jiang et al., 2024a), and use Llama-Guard-3-8B (Llama Team, 2024) as the automatic safety judge.
Figure 7 reports the proportion of harmful answers under the two settings across Qwen3 variants. A consistent pattern emerges across model scales: enabling the thinking mode substantially increases the likelihood of producing harmful answers. For all three Qwen3 variants, turning on thinking leads to a higher harmful-answer rate, with increases from 35.4% to 41.4% (8B), 32.3% to 35.6% (14B), and 30.9% to 41.5% (32B).
Notably, this effect does not diminish with model scaling. In fact, the largest model exhibits the most pronounced gap, suggesting that stronger reasoning capacity alone does not guarantee safer behavior. Instead, explicit reasoning may create additional opportunities for models to rationalize or justify unsafe responses. Collectively, these results offer preliminary evidence that explicit reasoning trajectories may amplify safety risks, motivating a finer-grained investigation of how unsafe behavior is instantiated within intermediate reasoning steps rather than being determined solely by the final output.
Figure 7: Harmful answer rate (%) of Qwen3 models on 2k harmful queries from WildJailbreak, comparing thinking mode enabled vs. disabled. Safety is automatically judged by Llama-Guard-3-8B. Enabling thinking consistently increases the proportion of harmful answers across model scales.
Appendix C Additional Experimental Results
C.1 Analysis: PCA Analysis of 8B and 14B Models
We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering. The observed differences are visualized in Figure 8 and 9, while quantitative results across all configurations are reported in Table 7.
Figure 8: PCA of the Qwen3-8B representation space.
Figure 9: PCA of the Qwen3-14B representation space.
@@ C.1 Analysis: PCA Analysis of 8B and 14B Models
Both Safety Recomposition (SafR) and Safety Backtrack (SafB) substantially improve model safety, as reflected by the centroid shifts in Figure 8 and Figure 9 and by the corresponding increases in Safety Distance reported in Table C.2. SafB reliably pushes the harmful cluster farther from the decision boundary while keeping the reasoning cluster relatively stable, indicating a balanced improvement in safety with limited impact on reasoning.
C.2 Preservation of Reasoning Token Usage
Table C.2 reports the average token length on GPQA-Diamond and AIME across training methods. Overall, SafR and SafB produce token counts comparable to the Base model, suggesting that their safety gains are not achieved by shortening generations. For example, on Qwen3-32B, the Base model averages 5150.81/12634.11 tokens (GPQA/AIME), while SafR remains close at 4830.90/12548.96 and SafB at 6452.00/12318.87. In contrast, SafePath shows a pronounced reduction to 3925.96/10200.49, with similar downward shifts on Qwen3-8B and Qwen3-14B. These results are consistent with the view that our methods improve safety while largely preserving the model’s intrinsic reasoning behavior.
Method GPQA-Diamond AIME
[…]
Base 7553.84 14895.49
@@ C.2 Preservation of Reasoning Token Usage
Table 13: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3 models across different training methods.
Method Harmful $\Delta$ Reasoning $\Delta$
[…]
Base 12.996 0.000 13.983 0.000
SafB 19.615 +6.619 12.800 –1.183
SafR 19.615 +6.619 12.800 –1.183
SafeChain 13.609 +0.613 5.324 –8.659
[…]
Base 15.540 0.000 9.054 0.000
SafB 19.142 +3.602 9.632 +0.578
SafR 18.952 +3.412 8.328 –0.726
SafeChain 12.590 –2.951 4.265 –4.789
[…]
Base 11.197 0.000 12.180 0.000
SafB 15.158 +3.961 10.409 –1.771
SafR 16.182 +4.985 11.504 –0.676
SafeChain 11.576 +0.379 6.921 –5.259
Table 14: Safety distance and relative changes ($\Delta$) for Qwen3 models under different safety training methods.
Model S-B S-RB W-JB JB-B GPQA AIME
@@
Table 15: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. S-B denotes Sorry-bench, S-RB denotes StrongREJECT, W-JB denotes WildJailbreak, and JB-B denotes JailBreak Bench. Lower is better for S-B, S-RB, W-JB, and JB-B; higher is better for GPQA and AIME.
C.3 Preservation of Reasoning Token Usage
To further contextualize the safety–reasoning trade-off, we contrast two representative paradigms: (i) reasoning-for-safety (or reasoning-as-defense), which improves robustness by explicitly injecting safety-aware reasoning signals (e.g., reflection or step-level self-assessment), and (ii) safety-for-reasoning, which aims to enhance safety while preserving the integrity of an LRM’s original reasoning process.
Table 15 compares our COG-based variants with a representative reasoning-as-defense baseline, R2D. While R2D achieves strong gains on safety-related metrics (SafB/SafR and WildJailbreak-style benchmarks), it exhibits a substantial drop in reasoning performance on GPQA and AIME, indicating that heavy safety-oriented reasoning control may interfere with general reasoning quality. In contrast, our SafB/SafR maintain markedly stronger reasoning performance (GPQA/AIME) while remaining competitive on safety benchmarks, suggesting that trajectory-level interventions can better preserve reasoning behavior under safety constraints. Nevertheless, the results also indicate that fully preserving reasoning integrity while maximizing safety remains challenging, motivating further study on how safety constraints reshape intermediate reasoning trajectories.
Appendix D Case Study: COT Structure Stability
To further demonstrate that our method maintains structural consistency with the original model’s reasoning patterns, we conduct a case study on a problem from AIME 2024. As shown in Figure 10, SafR exhibits a highly consistent reasoning structure with Qwen3-8B, while Star-1 employs a fundamentally different approach and ultimately produces an incorrect answer.
Specifically, as highlighted in blue, both Qwen3-8B and SafR decompose the problem into two symmetric cases based on the starting direction, apply the same combinatorial formula $C(7,2)\times C(7,1)=147$ for each case, and correctly aggregate the results as $147+147=294$. In contrast, Star-1 (red highlights) introduces a "four types" classification strategy, leading to systematic double-counting and yielding $588=2\times 294$. This case illustrates that SafR preserves not only the base model’s reasoning steps but also its mathematical correctness.
Appendix E Example of Safety Failure in LRMs
E.1 Harm Misidentification
The example of Warning in Harm Misidentification is presented in Figure 14, where LRM answers the harmful quesiton as the regular one.
E.2 Self-Jailbreak: Benign Reframing
@@
E.3 Self-Jailbreak: Logical Fallacies
@@ E.4 Self-Jailbreak: Warning
Appendix F Prompt Design
@@ F.2 Classification Prompt
F.3 Safety Recomposition Prompt
For the Safety Recomposition stage, our prompt is constructed by concatenating several components. Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts" field in a JSON file. Then, we combine the "main_prompt", the selected "sub_prompt", and a "format_prompt" to form the final prompt. The detailed structure is as shown in Figure 18
F.4 Safety Backtrack Prompt
In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process. In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases", selected based on the classification result. This helps guide the model more smoothly and maintain coherence in the final output. The detailed structure is as shown in Figure 19
F.5 Integration Prompt
@@
Figure 10: Reasoning pattern comparison on an AIME 2024 problem (ground truth: 294). Blue highlights indicate where SafR and Qwen3-8B share identical reasoning structures. Red highlights show Star-1’s divergent approach leading to an incorrect answer (588).